# Beware of Calibration Data for Pruning Large Language Models

### v1

@@ Abstract
As large language models (LLMs) are widely applied across various fields, model compression has become increasingly crucial for reducing costs and improving inference efficiency. Post-training pruning is a promising method that does not require resource-intensive iterative training and only needs a small amount of calibration data to assess the importance of parameters. Previous research has primarily focused on designing advanced pruning methods, while different calibration data’s impact on pruning performance still lacks systematical exploration. We fill this blank and surprisingly observe that the effects of calibration data even value more than designing advanced pruning strategies, especially for high sparsity. Our preliminary exploration also discloses that using calibration data similar to the training data can yield better performance. As pre-training data is usually inaccessible for advanced LLMs, we further provide a self-generating calibration data synthesis strategy to construct feasible calibration data. We conduct experiments on the recent strong open-source LLMs (e.g., DCLM, and LLaMA-3), and the results show that the proposed method outperforms commonly used calibration data and can effectively enhance strong pruning methods (e.g., Wanda, OWL).
@@ 1 Introduction
Recently, Large Language Models (LLMs) have exhibited remarkable performance and enormous potential in Natural Language Processing (NLP) and Artificial Intelligence (AI) (OpenAI, 2022; 2023; Bubeck et al., 2023; Yang et al., 2023). The success of LLMs is closely tied to scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022): training language models with more parameters, using more data and greater computational resources leads to more powerful capabilities. However, LLMs with more parameters increase the difficulty and cost of deployment and inference. Therefore, much work has been devoted to compressing LLMs to achieve a trade-off between efficiency and performance, such as pruning (Frantar & Alistarh, 2023; Ma et al., 2023; Xia et al., 2024) and quantization (Frantar et al., 2023; Huang et al., 2024; Shao et al., 2024).
@@
The challenge of post-training pruning is how to perform training-free parameter importance estimation. Frantar & Alistarh (2023) note that simple parameter magnitude-based metrics perform poorly in post-training pruning with over 20% sparsity. Therefore, they use a small amount of calibration data to compute the inverse Hessian matrix, estimating parameter importance through second-order gradient information. Sun et al. (2024) propose a simpler method by using the product of weight magnitudes and the L2 norm of the corresponding input activations. Dong et al. (2024) utilize the genetic algorithm to search for the optimal combination of information from magnitude, activation, and gradient as an importance metric. Overall, current advanced parameter importance metrics rely on calibration data. Although most papers claim their pruning methods are robust to calibration data, Williams & Aletras (2024)’s empirical study challenges this view. They demonstrate the performance differences of various methods using different calibration data. Furthermore, our experiments revealed that the performance gains from selecting better calibration data can even surpass those of advanced pruning methods (Figure 1). Therefore, it is time to focus more research on calibration data.
However, many open questions regarding calibration data remain under-explored. For example, how does the impact of calibration data change with increased sparsity and structure of pruning? Can increasing the amount of calibration data narrow the performance gap between various datasets? What type of data is suitable for calibration? How do we select the appropriate calibration data in practice? In this paper, we investigate these questions. Our empirical results demonstrate that as sparsity and structure increase, the performance differences among different calibration data become more pronounced, and simply increasing the data volume does not reduce this disparity. We further find that the selection of calibration data is closely related to the LLM’s training data, with calibration data similar to the training data yielding better performance. Based on this, we propose two strategies, detection and self-generation, aimed at sampling appropriate calibration data for pruning in practical settings with unavailable training data. To evaluate the effectiveness of our proposed calibration data sampling method, we conduct experiments on DCLM, LLaMA-2, and LLaMA-3 models. The results show that our proposed method performs better than the commonly used calibration data and is compatible with strong pruning methods by substantially improving their performance.
@@ 2 Background
Model compression is a crucial way to improve inference efficiency by reducing the required memory, including pruning (Frantar & Alistarh, 2023; Sun et al., 2024; Ma et al., 2023; Yin et al., 2024a; Guo et al., 2023; Zhang et al., 2024b; Xia et al., 2024), quantization (Frantar et al., 2023; Xiao et al., 2023; Lin et al., 2024; Huang et al., 2024; Shao et al., 2024), low-rank decomposition (Kaushal et al., 2023; Yuan et al., 2024; Wang et al., 2024; Ji et al., 2024), etc. The enormous memory requirements and inefficient inference speeds for LLMs urgently necessitate model compression. However, many successful model compression methods have required substantial computational resources for retraining, which limits their application for LLMs in low-resource settings. Therefore, post-training compression, which does not require retraining, has become a current research focus.
@@
where , are the original and compressed -th linear layer, respectively, and is the input feature activation. For post-training pruning, to optimize the objective, Frantar & Alistarh (2022; 2023) utilize second-order gradient information to measure parameter importance and propose an efficient algorithm for computing the inverse Hessian matrix. Sun et al. (2024) evaluate weight importance simply by the product of weight magnitudes and the norm of the corresponding input activation without requiring backpropagation. Zhang et al. (2024c) propose the relative importance and activation metric, which integrates weight, input, and output activation. They also utilize the channel permutation to minimize pruning loss under N:M semi-structured pruning. Dong et al. (2024) propose a search framework that employs the genetic algorithm to discover the optimal pruning metric for LLMs automatically. Recently, several studies (Sung et al., 2024; Xu et al., 2024a; Yin et al., 2024b) indicate that layer-wise compression, which typically applies a uniform sparsity rate across all layers and evaluates weight importance within the layer, often results in suboptimal performance due to the lack of overall consideration. Specifically, Xu et al. (2024a) proposes a differentiable pruning framework designed to search for optimal pruning rates for each layer. Yin et al. (2024b) introduce outlier weighed layerwise sparsity, which relates the sparsity of each layer to the observed outliers in a proportional manner.
@@
In the aforementioned post-training compression methods, calibration data is an indispensable component. Calibration data is a small subset randomly sampled from unlabeled pretraining text. Many methods (Frantar & Alistarh, 2023; Sun et al., 2024; Dettmers et al., 2024) claim their robustness to the quantity and distribution of calibration data, requiring only dozens or hundreds of samples with 2,048 sequence length. However, this conclusion is based on the perplexity of certain datasets (such as Wikitext2), which does not fully reflect the true capabilities of the LLMs. Even if perplexity shows no significant change, the compressed model may still experience substantial performance declines in downstream tasks (Jaiswal et al., 2024). Williams & Aletras (2024) observe in extensive experiments that the selection of calibration data in post-training pruning and quantization methods significantly impacts downstream tasks’ performance, especially post-training pruning, which is highly sensitive to calibration data. Nevertheless, current research on calibration data remains under-explored, with few studies providing guidelines for selecting calibration data. Khanal & Capone (2024) suggest that using task-specific calibration data helps improve performance on specific downstream tasks. Unlike their research, this paper aims to provide guidance on selecting calibration data to enhance the general capabilities of compressed models.
@@ 3.1 Experimental details
To study the impact of data from different sources on post-training pruning methods, we need a comprehensive knowledge of the data used in model training. We select the powerful and fully open-source LLM (including training data), DCLM-7B111https://huggingface.co/apple/DCLM-7B (Li et al., 2024), as the dense model and conduct post-training pruning with different calibration data on it.
@@
C4 (Raffel et al., 2020)222https://huggingface.co/datasets/allenai/c4 is a widely used calibration data source, consisting of a large amount of multilingual web text filtered from Common Crawl. We sample from the English training set.
Wikipedia333https://huggingface.co/datasets/wikimedia/wikipedia is a source of high-quality encyclopedic text. We use the first shard of the cleaned English version until 2023-11-01.
Slimpajama444https://huggingface.co/datasets/DKYoon/SlimPajama-6B is a cleaned and deduplicated version of RedPajama. It is a high-quality pre-training corpus with diverse sources, including C4, ArXiv, GitHub, Books, etc.
DCLM (Li et al., 2024) is the pre-training data of DCLM-7B model. It includes 2.6T tokens extracted from Common Crawl. We sample from a subset555https://huggingface.co/datasets/robbiegwaldd/dclm-micro of the DCLM.
@@
Some pruning works focus on the perplexity of certain datasets while neglecting performance on various downstream tasks, which often fails to fully reflect the capabilities of compressed models. Therefore, we choose multiple widely used and challenging commonsense reasoning tasks for evaluation, including BoolQ (Clark et al., 2019), Winogrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), Hellaswag (Zellers et al., 2019), ARC-e, ARC-c (Clark et al., 2018) and MMLU (Hendrycks et al., 2021). For MMLU, we use a 5-shot setting, while all other tasks are evaluated in a zero-shot setting. Our evaluation code is based on the lm-evaluation-harness repository666https://github.com/EleutherAI/lm-evaluation-harness. We report the average performance of these seven tasks.
@@ 3.4 What Calibration Data is Suitable for Pruning?
To verify the hypotheses, we perform three post-training pruning methods on DCLM-7B with various calibration data in the 2:4 semi-structured pruning setting. We report our results in Table 1.
@@
Among these data, using DCLM from the training data as calibration data consistently achieves the best performance. C4 and Slimpajama, which are also extracted from Common Crawl, perform slightly worse. In contrast, the source of Wikipedia differs significantly from the other three datasets. Although Wikipedia is recognized as high-quality data, it shows the worst performance, falling short of DCLM by 1.3% to 1.8%. Therefore, we assert that the quality of calibration data is not the primary factor affecting pruning performance. We further quantify the similarity between different calibration data and the training data. We utilize the MinHash-LSH algorithm to encode the 3-grams of C4, SlimPajama, Wikipedia, and DCLM, calculating their Jaccard similarities. The results show that the Jaccard similarity between C4 and DCLM is 0.070, SlimPajama is 0.016, and Wikipedia is 0.008. This indicates that C4 is the most similar to the training data, followed by SlimPajama, while Wikipedia has the lowest similarity. This ranking aligns with their performance as calibration data in pruning. Therefore, we believe that the similarity of calibration data to the training data has a more significant impact on pruning performance than the quality of the calibration data. Training data or data similar to the training data is better suited as calibration data. We conjecture that this may be due to LLMs learning the patterns in the training data better. Therefore, using data with similar patterns as calibration data during the pruning process can more accurately reflect the importance of model parameters.
@@ 5.1 Experimental Details
When replicating DSnoT and OWL, we follow the hyperparameter settings detailed in their papers. During the self-generation process, we use Top- and Top- sampling to improve the diversity of the generated data. Specifically, we set the -value to 0.95, the -value to 50, and the temperature to 0.6 for the DCLM model, and 0.8 temperature for the LLaMA-series model. We apply the repetition penalty of 1.2 to avoid repeatedly generating low-quality fragments. To demonstrate the generalization of our self-generated calibration data, we randomly sample 5,000 examples from the Wikipedia data for generation, as Wikipedia performs poorly among the LLMs we used. In the filtering phase, we eliminate the top 20% of samples based on their perplexity.
@@ 5.2 Overall Performance
| Method | Data | Alpaca () | BoolQ | Winogrande | PIQA | Hellaswag | ARC-e | ARC-c | MMLU | Avg. |
| DCLM-7B | ||||||||||
| Wanda | C4 | 9.67 | 78.47 | 70.27 | 75.12 | 66.32 | 72.84 | 40.84 | 43.31 | 63.88 |
| Wiki | 9.99 | 72.05 | 68.40 | 74.33 | 64.79 | 73.14 | 39.91 | 42.20 | 62.12 | |
| Slim | 9.76 | 78.56 | 70.16 | 74.27 | 65.07 | 72.37 | 39.94 | 43.40 | 63.40 | |
| DCLM | 9.54 | 79.11 | 70.51 | 75.13 | 66.25 | 73.37 | 41.66 | 44.58 | 64.37 | |
| Syn | 9.40 | 78.73 | 70.06 | 75.78 | 66.16 | 74.34 | 42.83 | 45.04 | 64.71 | |
| DSnoT | C4 | 9.81 | 76.11 | 69.44 | 74.76 | 65.08 | 72.10 | 39.08 | 41.62 | 62.60 |
| Wiki | 10.16 | 69.97 | 68.08 | 73.95 | 63.23 | 72.09 | 38.69 | 41.63 | 61.09 | |
| Slim | 9.87 | 75.58 | 69.21 | 73.80 | 63.88 | 71.37 | 38.63 | 42.25 | 62.10 | |
| DCLM | 9.70 | 77.39 | 69.36 | 74.63 | 64.89 | 72.06 | 39.83 | 43.73 | 63.13 | |
| Syn | 9.40 | 77.58 | 69.20 | 75.38 | 64.76 | 73.27 | 41.66 | 44.53 | 63.77 | |
| OWL | C4 | 9.52 | 78.14 | 68.90 | 75.55 | 65.22 | 72.46 | 38.24 | 39.04 | 62.51 |
| Wiki | 9.96 | 75.27 | 67.11 | 74.25 | 63.07 | 73.01 | 38.35 | 38.75 | 61.40 | |
| Slim | 9.59 | 78.09 | 68.69 | 74.56 | 64.00 | 72.35 | 37.95 | 39.84 | 62.21 | |
| DCLM | 9.38 | 78.45 | 69.47 | 75.10 | 65.07 | 72.76 | 38.81 | 40.73 | 62.91 | |
| Syn | 9.20 | 78.45 | 68.92 | 76.03 | 65.18 | 73.72 | 40.29 | 42.73 | 63.61 | |
| LLaMA-2-7B | ||||||||||
| Wanda | C4 | 10.42 | 66.30 | 64.50 | 71.12 | 58.92 | 64.92 | 33.91 | 23.06 | 54.68 |
| Wikipedia | 10.42 | 66.80 | 63.84 | 70.55 | 56.69 | 64.78 | 34.23 | 22.94 | 54.26 | |
| Slimpajama | 10.23 | 66.83 | 63.68 | 71.10 | 57.54 | 64.68 | 33.98 | 22.95 | 54.39 | |
| DCLM | 9.88 | 68.92 | 64.25 | 71.15 | 58.72 | 64.81 | 33.98 | 23.65 | 55.07 | |
| Syn | 9.62 | 68.29 | 64.40 | 71.49 | 58.89 | 64.73 | 35.41 | 24.01 | 55.32 | |
| DSnoT | C4 | 10.88 | 65.25 | 64.04 | 71.22 | 57.15 | 64.40 | 32.82 | 23.45 | 54.05 |
| Wikipedia | 10.92 | 66.24 | 62.72 | 70.55 | 55.55 | 64.10 | 33.16 | 23.05 | 53.62 | |
| Slimpajama | 10.76 | 65.66 | 63.66 | 70.82 | 56.17 | 64.43 | 32.51 | 23.15 | 53.77 | |
| DCLM | 10.37 | 66.65 | 63.99 | 71.44 | 56.77 | 64.56 | 33.30 | 23.73 | 54.35 | |
| Syn | 10.40 | 65.44 | 64.01 | 71.49 | 57.77 | 64.86 | 34.30 | 23.90 | 54.54 | |
| OWL | C4 | 9.19 | 66.73 | 67.34 | 72.74 | 62.86 | 67.54 | 35.68 | 26.20 | 57.02 |
| Wikipedia | 9.30 | 66.50 | 66.05 | 71.82 | 61.90 | 67.57 | 35.89 | 26.07 | 56.54 | |
| Slimpajama | 9.21 | 67.52 | 66.91 | 72.32 | 62.25 | 66.70 | 34.91 | 26.05 | 56.67 | |
| DCLM | 9.08 | 69.79 | 67.94 | 72.39 | 62.73 | 67.06 | 35.85 | 26.45 | 57.46 | |
| Syn | 9.13 | 69.85 | 66.38 | 73.18 | 62.86 | 67.89 | 35.07 | 26.34 | 57.37 | |
@@
We report main results in Table 2 and Table 5. Overall, our self-generated synthetic calibration data outperforms other baseline calibration data in language modeling and commonsense reasoning tasks and is compatible with different pruning methods. On DCLM-7B, the synthetic calibration data improves performance in commonsense reasoning tasks by an average of 2.2% to 2.6% compared to the original Wikipedia data. Additionally, it surpasses the commonly used C4 calibration data, achieving an average increase of 0.8% to 1.2%. On LLaMA family models, the self-generated synthetic data significantly outperforms the original data, with improvements ranging from approximately 0.9% to 1.1%, and surpasses the C4 data by about 0.3% to 0.5%. Surprisingly, the performance of the self-generated calibration data even exceeds that of calibration data sampled from the DCLM-7B training set, with an average improvement of 0.3% to 0.7%. We think this may be due to certain patterns in the calibration data that LLMs have not adequately learned. Using these patterns as calibration data may misestimate the importance of parameters. In contrast, due to the nature of maximum likelihood training, self-generated calibration data typically generates patterns that LLMs have better learned, thus avoiding using underrepresented patterns as calibration data.

### v2

@@ Abstract
As large language models (LLMs) are widely applied across various fields, model compression has become increasingly crucial for reducing costs and improving inference efficiency. Post-training pruning is a promising method that does not require resource-intensive iterative training and only needs a small amount of calibration data to assess the importance of parameters. Recent research has enhanced post-training pruning from different aspects but few of them systematically explore the effects of calibration data, and it is unclear if there exist better calibration data construction strategies. We fill this blank and surprisingly observe that calibration data is also crucial to post-training pruning, especially for high sparsity. Through controlled experiments on important influence factors of calibration data, including the pruning settings, the amount of data, and its similarity with pre-training data, we observe that a small size of data is adequate, and more similar data to its pre-training stage can yield better performance. As pre-training data is usually inaccessible for advanced LLMs, we further provide a self-generating calibration data synthesis strategy to construct feasible calibration data. Experimental results on recent strong open-source LLMs (e.g., DCLM, and LLaMA-3) show that the proposed strategy can enhance the performance of strong pruning methods (e.g., Wanda, DSnoT, OWL) by a large margin (up to 2.68%)111Code is available at https://github.com/Dereck0602/calibration_data.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have exhibited remarkable performance and enormous potential in Natural Language Processing (NLP) and Artificial Intelligence (AI) (OpenAI, 2022; 2023; Bubeck et al., 2023; Yang et al., 2023). The success of LLMs is closely tied to scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022): training language models with more parameters, using more data and greater computational resources leads to more powerful capabilities. However, LLMs with more parameters increase the difficulty and cost of deployment and inference. Therefore, much work has been devoted to compressing LLMs to achieve a trade-off between efficiency and performance, such as pruning (Ma et al., 2023; Xia et al., 2024) and quantization (Frantar et al., 2023; Huang et al., 2024; Shao et al., 2024).
@@
The challenge of post-training pruning is how to perform training-free parameter importance estimation. Frantar & Alistarh (2023) note that simple parameter magnitude-based metrics perform poorly in post-training pruning with over 20% sparsity. Therefore, they use a small amount of calibration data to compute the inverse Hessian matrix, estimating parameter importance through second-order gradient information. Sun et al. (2024) propose a simpler method by using the product of weight magnitudes and the L2 norm of the corresponding input activations. Dong et al. (2024) utilize the genetic algorithm to search for the optimal combination of information from magnitude, activation, and gradient as an importance metric. Overall, current advanced parameter importance metrics rely on calibration data. Although most papers claim their pruning methods are robust to calibration data, Williams & Aletras (2024)’s empirical study challenges this view. They demonstrate the performance differences of various methods using different calibration data. Our experiments further revealed that the performance gains from selecting better calibration data can even surpass those of advanced pruning methods (Figure 1).
To learn more about calibration data, we design experiments to explore (1) the impact of calibration data with increased sparsity and varied pruning types, (2) the influence of the amount of calibration data, and (3) the selection strategy of calibration data. Our empirical results demonstrate that as sparsity increases, the performance differences among different calibration data become more pronounced, and simply increasing the data volume does not reduce this disparity. We further find that calibration data similar to the pretraining data yields better performance. Based on this, we propose the self-generation strategy to construct appropriate calibration data for pruning in practical settings with unavailable training data. To evaluate the effectiveness of our proposed calibration data sampling method, we conduct experiments on DCLM, LLaMA-2, and LLaMA-3 models. The results show that our proposed method performs better than the commonly used calibration data and is compatible with strong pruning methods by substantially improving their performance.
@@ 2 Background
Model compression is a crucial way to improve inference efficiency by reducing the required memory, including pruning (Guo et al., 2023; Zhang et al., 2024b; Xia et al., 2024), quantization (Xiao et al., 2023; Lin et al., 2024), low-rank decomposition (Kaushal et al., 2023; Yuan et al., 2024; Wang et al., 2024; Ji et al., 2024), etc. The enormous memory requirements and inefficient inference speeds for LLMs urgently necessitate model compression. However, many successful model compression methods have required substantial computational resources for retraining, which limits their application for LLMs in low-resource settings. Therefore, post-training compression, which does not require retraining, has become a current research focus.
@@
where , are the original and compressed -th linear layer, respectively, and is the input feature activation. For post-training pruning, to optimize the objective, OBC (Frantar & Alistarh, 2022) and SparseGPT (Frantar & Alistarh, 2023) utilize second-order gradient information to measure parameter importance and propose an efficient algorithm for computing the inverse Hessian matrix. Wanda (Sun et al., 2024) evaluates weight importance by combining their magnitudes with input activations without requiring backpropagation. Zhang et al. (2024c) propose the relative importance and activation metric (RIA), which integrates weight, input, and output activation. They also utilize the channel permutation to minimize pruning loss under N:M semi-structured pruning. Pruner-Zero (Dong et al., 2024) designs a genetic algorithm-based framework to automatically search the best pruning metric. Recently, several studies (Sung et al., 2024; Xu et al., 2024a; Yin et al., 2024) indicate that layer-wise compression, which typically applies a uniform sparsity rate across all layers and evaluates weight importance within the layer, often results in suboptimal performance due to the lack of overall consideration. Specifically, Xu et al. (2024a) proposes a differentiable pruning framework designed to search for optimal pruning rates for each layer. OWL (Yin et al., 2024) introduces outlier weighed layerwise sparsity, which relates the sparsity of each layer to the observed outliers in a proportional manner.
@@
In the aforementioned post-training compression methods, calibration data is an indispensable component. Calibration data is a small subset randomly sampled from unlabeled pretraining text. Many methods (Frantar & Alistarh, 2023; Sun et al., 2024; Dettmers et al., 2024) claim their robustness to the quantity and distribution of calibration data, requiring only dozens or hundreds of samples with 2,048 sequence length. However, this conclusion is based on the perplexity of certain datasets (such as Wikitext2), which does not fully reflect the true capabilities of the LLMs. Even if perplexity shows no significant change, the compressed model may still experience substantial performance declines in downstream tasks (Jaiswal et al., 2024). Khanal & Capone (2024) suggest that using task-specific calibration data helps improve performance on specific downstream tasks. Williams & Aletras (2024) observe in extensive experiments that the selection of calibration data in post-training pruning and quantization methods significantly impacts downstream tasks’ performance, especially post-training pruning, which is highly sensitive to calibration data. Shin et al. (2024) notice that the reconstruction error objective (Eq. 1) leads to overfitting on calibration data, and that self-generated calibration data can effectively mitigate the overfitting. Nevertheless, current research on calibration data remains under-explored, with few studies providing guidelines for selecting calibration data. Different from previous works, our paper (1) explores the impact of calibration data under varying sparsity ratios and types, (2) investigates the effect of data amount on various calibration data, not limited to the widely used C4 calibration data, (3) further addresses which calibration data is suitable for LLM pruning and provides a practical and effective method.
@@ 3.1 Experimental details
To study the impact of data from different sources on post-training pruning methods, we need a comprehensive knowledge of the data used in model training. We select the powerful and fully open-source LLM (including training data), DCLM-7B222https://huggingface.co/apple/DCLM-7B (Li et al., 2024), as the dense model and conduct post-training pruning with different calibration data on it.
@@
C4 (Raffel et al., 2020)333https://huggingface.co/datasets/allenai/c4 is a widely used calibration data source, consisting of a large amount of multilingual web text filtered from Common Crawl. We sample from the English training set.
Wikipedia444https://huggingface.co/datasets/wikimedia/wikipedia is a source of high-quality encyclopedic text. We use the first shard of the cleaned English version until 2023-11-01.
Slimpajama555https://huggingface.co/datasets/DKYoon/SlimPajama-6B is a cleaned and deduplicated version of RedPajama. It is a high-quality pre-training corpus with diverse sources, including C4, ArXiv, GitHub, Books, etc.
DCLM (Li et al., 2024) is the pre-training data of DCLM-7B model. It includes 2.6T tokens extracted from Common Crawl. We sample from a subset666https://huggingface.co/datasets/robbiegwaldd/dclm-micro of the DCLM.
@@
Some pruning works focus on the perplexity of certain datasets while neglecting performance on various downstream tasks, which often fails to fully reflect the capabilities of compressed models. Therefore, we choose multiple widely used and challenging commonsense reasoning tasks for evaluation, including BoolQ (Clark et al., 2019), Winogrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), Hellaswag (Zellers et al., 2019), ARC-e, ARC-c (Clark et al., 2018) and MMLU (Hendrycks et al., 2021). For MMLU, we use a 5-shot setting, while all other tasks are evaluated in a zero-shot setting. Our evaluation code is based on the lm-evaluation-harness repository777https://github.com/EleutherAI/lm-evaluation-harness. We report the average performance of these seven tasks.
@@ 3.4 What Calibration Data is Suitable for Pruning?
To verify the hypotheses, we perform three post-training pruning methods on DCLM-7B with various calibration data in the 2:4 semi-structured pruning setting. We report our results in Table 1. Among these data, using DCLM from the training data as calibration data consistently achieves the best performance. C4 and Slimpajama, which are also extracted from Common Crawl, perform slightly worse. In contrast, the source of Wikipedia differs significantly from the other three datasets. Although Wikipedia is recognized as high-quality data, it shows the worst performance, falling short of DCLM by 1.3% to 1.8%. Therefore, we assert that the quality of calibration data is not the primary factor affecting pruning performance. We further quantify the similarity between different calibration data and the training data. We utilize the MinHash-LSH algorithm to encode the 3-grams of C4, SlimPajama, Wikipedia, and DCLM, calculating their Jaccard similarities. The results show that the Jaccard similarity between C4 and DCLM is 0.070, SlimPajama is 0.016, and Wikipedia is 0.008. This indicates that C4 is the most similar to the training data, followed by SlimPajama, while Wikipedia has the lowest similarity. This ranking aligns with their performance as calibration data in pruning. Therefore, we believe that the similarity of calibration data to the training data has a more significant impact on pruning performance than the quality of the calibration data. Training data or data similar to the training data is better suited as calibration data. We conjecture that this may be due to LLMs learning the patterns in the training data better. Therefore, using data with similar patterns as calibration data during the pruning process can more accurately reflect the importance of model parameters.
@@ 5.1 Experimental Details
When replicating DSnoT and OWL, we follow the hyperparameter settings detailed in their papers. During the self-generation process, we use Top- and Top- sampling to improve the diversity of the generated data. Specifically, we set the -value to 0.95, the -value to 50, and the temperature to 0.6. We apply the repetition penalty of 1.2 to avoid repeatedly generating low-quality fragments. We randomly sample 5,000 examples from C4, Slimpajama, Wikipedia, and DCLM respectively for generation. In the filtering phase, we eliminate the top 20% of samples based on their perplexity.
@@ 5.2 Overall Performance
Data | Pretrain | Alpaca () | BoolQ | Winogrande | PIQA | Hellaswag | ARC-e | ARC-c | MMLU | Avg. | |
Wanda | |||||||||||
Wiki | ✗ | 9.99 | 72.05 | 68.40 | 74.33 | 64.79 | 73.14 | 39.91 | 42.20 | 62.12 | |
w/ Syn | 9.40 | 78.73 | 70.06 | 75.78 | 66.16 | 74.34 | 42.83 | 45.04 | 64.71 | +2.59 | |
C4 | ✓✗ | 9.67 | 78.47 | 70.27 | 75.12 | 66.32 | 72.84 | 40.84 | 43.31 | 63.88 | |
w/ Syn | 9.57 | 78.81 | 70.52 | 75.95 | 66.35 | 74.23 | 42.01 | 45.64 | 64.78 | +0.90 | |
Slim | ✓✗ | 9.76 | 78.56 | 70.16 | 74.27 | 65.07 | 72.37 | 39.94 | 43.40 | 63.40 | |
w/ Syn | 9.58 | 78.51 | 70.02 | 75.63 | 65.90 | 74.12 | 42.13 | 45.26 | 64.51 | +1.11 | |
DCLM | ✓ | 9.54 | 79.11 | 70.51 | 75.13 | 66.25 | 73.37 | 41.66 | 44.58 | 64.37 | |
w/ Syn | 9.59 | 79.23 | 70.69 | 75.64 | 66.17 | 74.04 | 42.01 | 45.42 | 64.74 | +0.37 | |
DSnoT | |||||||||||
Wiki | ✗ | 10.16 | 69.97 | 68.08 | 73.95 | 63.23 | 72.09 | 38.69 | 41.63 | 61.09 | |
w/ Syn | 9.40 | 77.58 | 69.20 | 75.38 | 64.76 | 73.27 | 41.66 | 44.53 | 63.77 | +2.68 | |
C4 | ✓✗ | 9.81 | 76.11 | 69.44 | 74.76 | 65.08 | 72.10 | 39.08 | 41.62 | 62.60 | |
w/ Syn | 9.56 | 75.61 | 69.30 | 75.56 | 65.13 | 73.06 | 41.11 | 45.24 | 63.57 | +0.97 | |
Slim | ✓✗ | 9.87 | 75.58 | 69.21 | 73.80 | 63.88 | 71.37 | 38.63 | 42.25 | 62.10 | |
w/ Syn | 9.62 | 76.08 | 69.27 | 75.09 | 64.57 | 73.16 | 40.97 | 44.57 | 63.39 | +1.29 | |
DCLM | ✓ | 9.70 | 77.39 | 69.36 | 74.63 | 64.89 | 72.06 | 39.83 | 43.73 | 63.13 | |
w/ Syn | 9.52 | 76.56 | 68.35 | 75.55 | 64.70 | 73.43 | 41.43 | 44.81 | 63.55 | +0.42 | |
OWL | |||||||||||
Wiki | ✗ | 9.96 | 75.27 | 67.11 | 74.25 | 63.07 | 73.01 | 38.35 | 38.75 | 61.40 | |
w/ Syn | 9.20 | 78.45 | 68.92 | 76.03 | 65.18 | 73.72 | 40.29 | 42.73 | 63.61 | +2.21 | |
C4 | ✓✗ | 9.52 | 78.14 | 68.90 | 75.55 | 65.22 | 72.46 | 38.24 | 39.04 | 62.51 | |
w/ Syn | 9.31 | 78.55 | 68.67 | 76.38 | 65.45 | 74.05 | 40.03 | 42.94 | 63.72 | +1.21 | |
Slim | ✓✗ | 9.59 | 78.09 | 68.69 | 74.56 | 64.00 | 72.35 | 37.95 | 39.84 | 62.21 | |
w/ Syn | 9.32 | 78.56 | 68.71 | 75.83 | 64.47 | 73.81 | 40.44 | 43.61 | 63.64 | +1.43 | |
DCLM | ✓ | 9.38 | 78.45 | 69.47 | 75.10 | 65.07 | 72.76 | 38.81 | 40.73 | 62.91 | |
w/ Syn | 9.28 | 78.80 | 67.77 | 75.90 | 64.77 | 73.84 | 40.56 | 43.67 | 63.61 | +0.70 |
@@
We report the main results in Table 2 and Table 5. Overall, our self-generated synthetic calibration data exceeds other baseline calibration data in language modeling and commonsense reasoning tasks and is compatible with different pruning methods. On DCLM-7B, Wikipedia, which is not part of the pretraining data, achieves the greatest performance improvement through self-generating synthetic data. It improves performance in commonsense reasoning tasks by an average of 2.2% to 2.6% compared to the original Wikipedia data, and even surpasses the commonly used C4 calibration data, achieving an average increase of 0.8% to 1.2%. For C4 and Slimpajama, which partially overlap with the pretraining data, the self-generation strategy also yields a 0.9-1.5% improvement. On LLaMA family models, the self-generated synthetic data also performs better than the original data, with improvements ranging from approximately 0.9% to 1.1%, and surpasses the C4 data by about 0.3% to 0.5%. Surprisingly, the performance of the self-generated calibration data even exceeds that of calibration data sampled from the DCLM-7B training set, with an average improvement of 0.3% to 0.7%. We think this may be due to certain patterns in the calibration data that LLMs have not adequately learned. Using these patterns as calibration data may misestimate the importance of parameters. In contrast, due to the nature of maximum likelihood training, self-generated calibration data typically generates patterns that LLMs have better learned, thus avoiding using underrepresented patterns as calibration data. Additionally, we observe that regardless of the source of synthetic data, the pruned models’ performances are similar. It indicates that self-generated calibration data is versatile, as it can generate suitable calibration data even when the available data is significantly different from the pretraining data.